Solve each equation for x. Factor the left-hand side
x² +4x + 4 =7

Answer by stanbon(75887)

Solve each equation for x. Factor the left-hand side
x² +4x + 4 =7
-----------------
(x^2)^2 = 7
(x+2)^2-(sqrt7)^2 = 0
Factor:
(x+2-srt7)(x+2+sqrt7)=0
x=-2+sqrt7 or x= -2-sqrt7
==================
Cheers,
Stan H.